# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v11

@@ Abstract.
We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting. For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove that on a non-empty open subset $\mathcal{U}\subset M$,
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
@@ 1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold [^1] As is customary in Riemannian geometry, all manifolds are assumed to be connected unless otherwise stated. endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
A subtle but crucial point is that the subspaces spanned by a minimal harmonic decomposition are uniquely determined only under topological restrictions on the factors. Consequently, our main result holds for admissible product manifolds (Definition 2.2), i.e.:
  • (A)
    both factors are compact oriented surfaces;  or
  • (B)
    $\dim M=4$ with a $(3,1)$ splitting.
For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies, on a non-empty open subset $\mathcal{U}\subset M$,
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
@@ 2.1. Intrinsic Definition of Mixed Cohomology Subspaces
We now address a subtle but crucial point: the subspaces spanned by the factors in a minimal harmonic decomposition are not uniquely determined in general. To obtain a well–defined geometric invariant we must impose conditions on the topology of the factors.
@@
A compact oriented product manifold $M=M_{1}\times M_{2}$ is called admissible if it satisfies one of the following:
  • (A)
    both factors $M_{1},M_{2}$ are compact oriented surfaces;
  • (B)
    $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$. [^2] This is not an additional condition, but holds automatically: $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\geq 1$ because $\omega]_{\mathrm{mixed}}\neq 0$ (not trivial mixed class), while $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\leq 1$ since $S^{1}$ is the unique compact connected oriented $1$-manifold, with $\dim H^{1}(S^{1})=1$ and $H^{2}(S^{1})=0$; hence equality holds.
@@
We treat the two admissible cases separately.
Case (A): $\dim M_{i}=2$. For a compact oriented surface $F^{2}$ we have $b_{2}(F)=1$ and $b_{1}(F)=2g(F)$. Let $\operatorname{vol}_{F}$ be the unique (up to scale) $g_{F}$–harmonic 22–form. Then
@@
is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$.
Case (B): $(3,1)$ or $(1,3)$. Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric). Then $b_{1}(M_{2})=1$ and $H^{2}(M_{2})=0$. Let $\gamma$ be a fixed basis of $\mathcal{H}^{1}(M_{2})$. Any $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank at most one and can be written uniquely as
$T=\alpha\otimes\gamma,\qquad\alpha\in\mathcal{H}^{2}(M_{1}).$
Uniqueness follows as in Case (A) because $\gamma$ is non–zero. Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ is intrinsically defined. The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined. ∎
@@ 2.2. Mixed obstruction spaces
The notation $\ker\Psi_{p}$” is anticipatory: in the proof of Theorem 5.2, we construct obstruction maps $\Psi_{p}:V_{2,1}\to\mathfrak{so}(T_{p}M)$ and $\widetilde{\Psi}_{p}:V_{1,2}\to\mathfrak{so}(T_{p}M)$ via covariant derivatives of the torsion form. A key step in the proof is to verify that the kernels of these maps coincide with the intersections defined above. The induced quotient maps $V_{2,1}/\ker\Psi_{p}\to\mathfrak{so}(T_{p}M)$ are then shown to be injective on a non-empty open subset, establishing the lower bound.
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
Although $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is a vector subspace of $\mathfrak{hol}_{p}(\nabla^{C})$, it is not closed under the Lie bracket, since $\mathfrak{so}^{\mathrm{off}},\mathfrak{so}^{\mathrm{off}}]\subset\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$. Nevertheless, $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})>0$ implies that the holonomy representation is irreducible with respect to the splitting $V_{1}\oplus V_{2}$.
@@ 5. Main Result and Proof
If $r-s\leq 1$ (condition satisfied by admissible product manifolds: $U$ is the span of a minimal harmonic decomposition, so $\dim U=r_{2,1}\leq 1$, hence $r-s\leq 1$; see Definition 4.1, Definition 2.2 and Lemma 2.3), then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
@@
If we prove that for each $k$ there exists a non‑empty open set $\mathcal{U}^{(k)}\subset M_{1}$ where the $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}$ are linearly independent, then the $C^{1}$ convergence guarantees that, for large enough $k$, the same linear independence holds on a non‑empty open set for the original metric $g_{1}$. Indeed, linear independence is equivalent to the non‑vanishing of at least one $(r-s)\times(r-s)$ minor of the matrix of components; each minor depends continuously on the metric and the forms, so a minor that is non‑zero on an open set for $g_{1}^{(k)}$ remains non‑zero on an open set for $g_{1}$ when $k$ is sufficiently large.
@@
Case $r=s+1$: There exists a unique non‑parallel form $\alpha$ and a single analytic function $c$ such that (5.1) reads $c(p)(\nabla^{g_{1}}\alpha)_{p}=0$ for all $p\in V$. If $c(p_{0})\neq 0$ for some $p_{0}\in V$, then by analyticity $c(p)\neq 0$ on an open subset $V^{\prime}\subset V$, whence $(\nabla^{g_{1}}\alpha)_{p}=0$ on $V^{\prime}$. By the unique‑continuation principle for analytic sections (Aronszajn [1]; see also [8]), $\nabla^{g_{1}}\alpha\equiv 0$ on the whole connected component containing $V^{\prime}$, and hence on all of $M_{1}$. This contradicts the assumption that $\alpha$ is non‑parallel modulo $\mathcal{P}_{2}(M_{1},g_{1})$. Therefore $c\equiv 0$ on $V$, contradicting the choice of the $c_{i}$’s. Hence $\mathcal{U}\neq\emptyset$.
3. Returning to the original smooth metric. As explained in Step 1, the existence of a non‑empty open set of linear independence for a sequence of real‑analytic approximations $g_{1}^{(k)}$ together with the $C^{1}$ convergence of the covariant derivatives guarantees the existence of a non‑empty open set with the same property for the original smooth metric $g_{1}$. This completes the proof. ∎
@@
Let $M=M_{1}\times M_{2}$ be an admissible compact oriented Riemannian product manifold (Definition 2.2) with product metric $g=g_{1}\oplus g_{2}$. Let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]$ is a non‑trivial mixed class.
@@
Then, there exists a non-empty open subset $\mathcal{U}\subset M$ such that for every $p\in\mathcal{U}$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
@@
Set $r^{\sharp}:=r-\dim\mathcal{K}$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
Consider the curvature formula (2.5):
$R^{C}=R^{LC}+\frac{1}{2}(\nabla^{LC}T)+\frac{1}{4}Q_{T}.$
The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts. Taking off-diagonal components:
$R^{C}_{\mathrm{off}}=R^{LC}_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\eta)_{\mathrm{off}}+\frac{1}{4}(Q_{T})_{\mathrm{off}}.$
Our aim is to show that $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}:=r-\dim\mathcal{K}$ on a non-empty open subset of $M$. We first establish that to prove this bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic term $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ alone.
Suppose that at some point $p\in M$ we have constructed operators $A_{1},\ldots,A_{r^{\sharp}}\in\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ that are linearly independent and obtained from evaluations of $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$. By the curvature formula, each such operator is a value of $R^{C}_{\mathrm{off}}$, since $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ appears as one of the terms in $R^{C}_{\mathrm{off}}$. Let
$S_{p}^{\mathrm{off}}:=\mathrm{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(R^{C}(X,Y)):X,Y\in T_{p}M\}$
be the span of all off-diagonal curvature values at $p$. By the curvature formula (2.5), the covariant derivative $(\nabla^{LC}_{X}T)(Y,Z,\cdot)$ appears in the operator $R^{C}(X,Y)$ after raising indices; hence the operators $A_{1},\ldots,A_{r^{\sharp}}$ constructed from evaluations of $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ are indeed components of curvature values, and therefore belong to $S_{p}^{\mathrm{off}}$. Since they are linearly independent, we have $\dim S_{p}^{\mathrm{off}}\geq r^{\sharp}$. The terms $R^{LC}_{\mathrm{off}}$, $(\nabla^{LC}\eta)_{\mathrm{off}}$, and $(Q_{T})_{\mathrm{off}}$ contribute additional values to $S_{p}^{\mathrm{off}}$, but cannot reduce its dimension below $r^{\sharp}$. By the Ambrose–Singer theorem [2], $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ contains $S_{p}^{\mathrm{off}}$ as the off-diagonal part of the algebra generated by curvature values, whence $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$. Thus, to establish the desired lower bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic component $\omega_{h}$.
@@
To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists a non‑empty open subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
@@
$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}),$ (5.4)
where the vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfy $\alpha_{i}(X_{i},Y_{i})=1$ and $\beta_{i}(Z_{i})=1$.
Claim 5.2.1 (Independence of choices). The subspace $\ker\Psi_{p}\subset\mathcal{V}_{2,1}$ is independent of the choice of normalizing vectors $\{X_{i},Y_{i},Z_{i}\}$ in  5.4.

### v12

@@ Abstract.
We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on product manifolds where both factors are compact oriented surfaces. For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove that on a non-empty open subset $\mathcal{V}\subset M$,
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\min\bigl(\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},\;1\bigr),$
@@ 1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
A subtle but crucial point is that the subspaces spanned by a minimal harmonic decomposition are uniquely determined only under topological restrictions on the factors. Consequently, our main result holds for admissible product manifolds (Definition 2.2), i.e., where both factors are compact oriented surfaces.
For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies, on a non-empty open subset $\mathcal{V}\subset M$,
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\min\bigl(\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},\;1\bigr),$
@@ 2.1. Intrinsic Definition of Mixed Cohomology Subspaces
We now address a subtle but fundamental point: the subspaces spanned by the factors in a minimal harmonic decomposition are not uniquely determined in general. To obtain a well–defined geometric invariant we must impose conditions on the topology of the factors.
@@
A compact oriented product manifold $M=M_{1}\times M_{2}$ is called admissible if it satisfies the following:
  • both factors $M_{1}$, $M_{2}$ are compact oriented surfaces.
@@
For a compact oriented surface $F^{2}$ we have $b_{2}(F)=1$ and $b_{1}(F)=2g(F)$. Let $\operatorname{vol}_{F}$ be the unique (up to scale) $g_{F}$–harmonic 22–form. Then
@@
is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$. ∎
@@ 2.2. Mixed obstruction spaces
The notation $\ker\Psi_{p}$” is anticipatory: in the proof of Theorem 5.2, we construct obstruction maps $\Psi_{p}:V_{2,1}\to\mathfrak{so}(T_{p}M)$ and $\widetilde{\Psi}_{p}:V_{1,2}\to\mathfrak{so}(T_{p}M)$ via covariant derivatives of the torsion form. A key step in the proof is to verify that the kernels of these maps coincide with the intersections defined above. The induced quotient maps $V_{2,1}/\ker\Psi_{p}\to\mathfrak{so}(T_{p}M)$ are then shown to be injective on a dense open subset, establishing the lower bound.
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
Although $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is a vector subspace of $\mathfrak{hol}_{p}(\nabla^{C})$, it is not closed under the Lie bracket, since $\mathfrak{so}^{\mathrm{off}},\mathfrak{so}^{\mathrm{off}}]\subset\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$. Nevertheless, $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})>0$ implies that the holonomy representation is irreducible with respect to the splitting $V_{1}\oplus V_{2}$, therefore, we have global irreducibility.
@@ 5. Main Result and Proof
If $r-s\leq 1$ (condition satisfied by admissible product manifolds: $U$ is the span of a minimal harmonic decomposition, so $\dim U=r_{2,1}\leq 1$, hence $r-s\leq 1$; see Definition 4.1, Definition 2.2 and Lemma 2.3), then there exists a dense open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
@@
If we prove that for each $k$ there exists a dense open set $\mathcal{U}^{(k)}\subset M_{1}$ where the $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}$ are linearly independent, then the $C^{1}$ convergence guarantees that, for large enough $k$, the same linear independence holds on a dense open set for the original metric $g_{1}$. Indeed, linear independence is equivalent to the non‑vanishing of at least one $(r-s)\times(r-s)$ minor of the matrix of components; each minor depends continuously on the metric and the forms, so a minor that is non‑zero on an open set for $g_{1}^{(k)}$ remains non‑zero on an open set for $g_{1}$ when $k$ is sufficiently large.
@@
Case $r=s+1$: There exists a unique non‑parallel form $\alpha$ and a single analytic function $c$ such that (5.1) reads $c(p)(\nabla^{g_{1}}\alpha)_{p}=0$ for all $p\in V$. If $c(p_{0})\neq 0$ for some $p_{0}\in V$, then by analyticity $c(p)\neq 0$ on an open subset $V^{\prime}\subset V$, whence $(\nabla^{g_{1}}\alpha)_{p}=0$ on $V^{\prime}$. By the unique‑continuation principle for analytic sections (Aronszajn [1]; for differential forms see [2]), $\nabla^{g_{1}}\alpha\equiv 0$ on the whole connected component containing $V^{\prime}$, and hence on all of $M_{1}$. This contradicts the assumption that $\alpha$ is non‑parallel modulo $\mathcal{P}_{2}(M_{1},g_{1})$. Therefore $c\equiv 0$ on $V$, contradicting the choice of the $c_{i}$’s. Hence $\mathcal{U}\neq\emptyset$. Moreover, the complement $M_{1}\setminus\mathcal{U}=\{p\in M_{1}:(\nabla^{g_{1}}\alpha)(p)=0\}$ has empty interior: if it contained an open set $V^{\prime}$, then by unique continuation $\nabla^{g_{1}}\alpha\equiv 0$ on all of $M_{1}$, contradicting the hypothesis that $\alpha$ is non-parallel. Therefore $\mathcal{U}$ is dense in $M_{1}$.
3. Returning to the original smooth metric. By Step 2, for each analytic approximation $g_{1}^{(k)}$, the set $\mathcal{U}^{(k)}=\{p:\nabla^{g_{1}^{(k)}}\alpha^{(k)}(p)\neq 0\}$ is dense in $M_{1}$. The same conclusion holds for the smooth metric $g_{1}$ by applying the unique continuation principle for harmonic forms on smooth Riemannian manifolds [2] directly: if $M_{1}\setminus\mathcal{U}$ contained an open set $V$, lifting to the universal cover $\tilde{M}_{1}$ yields a globally parallel form $\tilde{\beta}$ with $\tilde{\alpha}|_{\tilde{V}}=\tilde{\beta}|_{\tilde{V}}$. Since $\tilde{\alpha}-\tilde{\beta}$ is harmonic and vanishes on $\tilde{V}$, unique continuation [2] implies $\tilde{\alpha}=\tilde{\beta}$, so $\alpha$ is parallel on $M_{1}$, contradicting $\alpha\notin U_{\parallel}$. Thus $\mathcal{U}$ is dense. ∎
@@
Let $M=M_{1}\times M_{2}$ be an admissible compact oriented Riemannian product manifold (Definition 2.2) with product metric $g=g_{1}\oplus g_{2}$. Let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]$ is a non‑trivial mixed class (Definition 2.1).
@@
Then, there exists a non-empty open subset $\mathcal{V}\subset M$ such that for every $p\in\mathcal{V}$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\;\geq\;r^{\sharp}:=\min\bigl(\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},\;1\bigr).$
@@
Set $r^{\sharp}:=\min(r-\dim\mathcal{K},\,1)$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
Consider the curvature formula (2.5) for $X\in V_{1}$ and $Y\in V_{2}$:
$R^{C}(X,Y)\;=\;R^{LC}(X,Y)\;+\;\tfrac{1}{2}\bigl[(\nabla^{LC}_{X}T)(Y,\cdot\,,\cdot)^{\sharp}-(\nabla^{LC}_{Y}T)(X,\cdot\,,\cdot)^{\sharp}\bigr]\;+\;\tfrac{1}{4}\,Q_{T}(X,Y).$
On a Riemannian product the mixed Riemann curvature vanishes: $R^{LC}(X,Y)=0$ whenever $X\in V_{1}$ and $Y\in V_{2}$ (since $\nabla^{LC}_{X}Y=0=[X,Y]$ for vector fields tangent to different factors). For the remaining three terms we exploit the rigid pointwise algebraic structure that the surface hypothesis imposes on mixed $3$-forms.
Since both factors are compact oriented surfaces, $\dim V_{i}=2$, and the exterior power $\Lambda^{2}V_{1}^{*}$ is one-dimensional, spanned by the area form $\operatorname{vol}_{V_{1}}$. Every $3$-form of bidegree $(2,1)$ on $M$ can therefore be written at each point $p$ as
$T^{2,1}_{p}\;=\;h(p)\;\operatorname{vol}_{V_{1}}\wedge\tau_{p},$ (5.4)
where $h\colon M\to\mathbb{R}$ is a smooth function and $\tau$ is a smooth section of $V_{2}^{*}$ (depending on the full point $p=(p_{1},p_{2})$). Likewise, every $(1,2)$-form satisfies $T^{1,2}_{p}=\widetilde{h}(p)\;\widetilde{\sigma}_{p}\wedge\operatorname{vol}_{V_{2}}$ with $\widetilde{\sigma}\in\Gamma(V_{1}^{*})$. These representations hold for the full torsion $T$, not merely for its harmonic projection $\omega_{h}$.
We denote by $J_{i}\colon V_{i}\to V_{i}$ the volume endomorphism of the oriented surface $M_{i}$, characterised by $g(J_{i}u,v)=\operatorname{vol}_{V_{i}}(u,v)$; it is a skew-symmetric isometry with $J_{i}^{2}=-\operatorname{Id}$, and in particular $J_{i}X\perp X$ for every nonzero $X\in V_{i}$. Let $\pi_{0}\colon\mathfrak{so}(T_{p}M)\to\mathfrak{g}_{0}:=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$ and $\pi_{1}\colon\mathfrak{so}(T_{p}M)\to\mathfrak{g}_{1}:=\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ denote the projections induced by the $\mathbb{Z}_{2}$-grading of $\mathfrak{so}(T_{p}M)$.
[…]
$\pi_{0}\bigl(R^{C}(X,Y)\bigr)\;=\;\lambda(X,Y)\,J_{1}\;+\;\mu(X,Y)\,J_{2}$ (5.5)
[…]
$\rho_{Y}\;:=\;(Yh)\,\tau\;+\;h\,\nabla^{g_{2}}_{Y}\!\tau\;=\;\nabla^{g_{2}}_{Y}(h\tau),$ (5.6)
the restriction to $\operatorname{Hom}(V_{1},V_{2})$ reads
$\tfrac{1}{2}(\nabla^{LC}_{Y}T^{2,1})(X,Z,\cdot)^{\sharp}\Big|_{Z\in V_{1}}\;=\;-\tfrac{1}{2}\,g(J_{1}X,Z)\;\rho_{Y}^{\sharp}\;\in V_{2}.$ (5.7)
The third, $\frac{1}{4}Q_{T}(X,Y)$, is also purely off-diagonal (by a bidegree computation), with restriction
$\tfrac{1}{4}\,Q_{T}(X,Y)\big|_{Z\in V_{1}}\;=\;\tfrac{h^{2}}{4}\;\tau(Y)\;g(X,Z)\;\tau^{\sharp}\;\in V_{2}.$ (5.8)
To derive 5.8, observe that the contortion tensor $K_{X}Y=\tfrac{1}{2}T(X,Y,\cdot)^{\sharp}$ acts on $Z\in V_{1}$ via the $(2,1)$-component as
$K_{Y}^{(2,1)}Z=\frac{h}{2}\,\tau(Y)\,J_{1}Z,$
since $(\operatorname{vol}_{V_{1}}\wedge\tau)(Y,Z,W)=\tau(Y)\,\operatorname{vol}_{V_{1}}(Z,W)$ for $Y\in V_{2}$ and $Z,W\in V_{1}$. The iterated contortion $K_{X}(K_{Y}Z)$ with $X\in V_{1}$ then gives
$K_{X}(K_{Y}Z)=\frac{h^{2}}{4}\,\tau(Y)\,g(X,Z)\,\tau^{\sharp},$
where we used $T^{(2,1)}(X,J_{1}Z,W)=h\,\operatorname{vol}_{V_{1}}(X,J_{1}Z)\,\tau(W)=h\,g(J_{1}X,J_{1}Z)\,\tau(W)=h\,g(X,Z)\,\tau(W)$, the last equality following from $J_{1}$ being an isometry. Conversely, $K_{Y}(K_{X}Z)=0$: since $K_{X}Z=\tfrac{h}{2}\operatorname{vol}_{V_{1}}(X,Z)\,\tau^{\sharp}\in V_{2}$ and $T^{(2,1)}(Y,W,\cdot)=0$ whenever $Y,W\in V_{2}$ (the form $\operatorname{vol}_{V_{1}}$ requires two arguments in $V_{1}$). Therefore $Q_{T}(X,Y)Z=K_{X}(K_{Y}Z)-K_{Y}(K_{X}Z)$ yields 5.8.
Combining (5.7) and (5.8), the full off-diagonal projection restricted to $\operatorname{Hom}(V_{1},V_{2})$ is
$\pi_{1}\bigl(R^{C}(X,Y)\bigr)\big|_{V_{1}\to V_{2}}\;:\;Z\;\longmapsto\;-\tfrac{1}{2}\,g(J_{1}X,Z)\;\rho_{Y}^{\sharp}\;+\;\tfrac{h^{2}}{4}\,\tau(Y)\,g(X,Z)\;\tau^{\sharp}.$ (5.9)
The two summands act through the functionals $g(J_{1}X,\cdot\,)$ and $g(X,\cdot\,)$ on $V_{1}$ respectively. Because $J_{1}X\perp X$, these functionals are orthogonal, and the two contributions cannot cancel: if the map (5.9) vanishes for all $Z\in V_{1}$, then
$\rho_{Y}^{\sharp}=0\qquad\text{and}\qquad h^{2}\,\tau(Y)\,\tau^{\sharp}=0$ (5.10)
must hold separately.
We now observe that the quadratic term (5.8) is non-vanishing on a non-empty open set. If $r^{\sharp}_{2,1}=1$, the harmonic projection $\omega_{h}^{2,1}=c\cdot\operatorname{vol}_{M_{1}}\wedge\beta$ is non-zero, so $T^{2,1}]=[\omega_{h}^{2,1}]\neq 0$ in cohomology. Since $T^{2,1}=\omega_{h}^{2,1}+\eta^{2,1}$ with $\eta^{2,1}$ exact plus coexact, Hodge orthogonality gives $T^{2,1}\not\equiv 0$ on $M$ (otherwise $\omega_{h}^{2,1}=-\eta^{2,1}$ would be simultaneously harmonic and exact-plus-coexact, forcing $\omega_{h}^{2,1}=0$). The set
$\mathcal{W}^{2,1}\;:=\;\{p\in M:T^{2,1}_{p}\neq 0\}$ (5.11)
is therefore a non-empty open subset of $M$. On $\mathcal{W}^{2,1}$ we have $h(p)\neq 0$ and $\tau_{p}\neq 0$ (by (5.4)), so choosing $Y\in V_{2}$ with $\tau_{p}(Y)\neq 0$ yields $h^{2}\tau(Y)\tau^{\sharp}\neq 0$; the second condition in (5.10) is therefore violated, and $\pi_{1}(R^{C}(X,Y))\neq 0$. The symmetric argument for the $(1,2)$-component (using $J_{2}Y\perp Y$) produces a non-empty open set $\mathcal{W}^{1,2}$ whenever $r^{\sharp}_{1,2}=1$.
In particular, the exact-plus-coexact part $\eta$ and the quadratic torsion term $Q_{T}$ cannot reduce the number of independent off-diagonal curvature directions below the topological invariant $r^{\sharp}$. To compute this invariant explicitly and to construct $r^{\sharp}$ independent off-diagonal operators, it therefore suffices to work with the harmonic projection $\omega_{h}$ alone, as we now proceed to do.
@@
To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists a dense open subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
@@
$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}\omega_{h})^{\mathrm{off}}(X_{i},Y_{i},Z_{i}),$ (5.12)
where the vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfy $\alpha_{i}(X_{i},Y_{i})=1$ and $\beta_{i}(Z_{i})=1$. The notation on the right-hand side denotes the off-diagonal component of the curvature contribution from the $\nabla^{LC}\omega_{h}$-term in 2.5, and its precise meaning is established through the kernel characterization in Claim 5.2.1.
Claim 5.2.1 (Independence of choices). The subspace $\ker\Psi_{p}\subset\mathcal{V}_{2,1}$ is independent of the choice of normalizing vectors $\{X_{i},Y_{i},Z_{i}\}$ in  5.12.